# Grand-Canonical Typicality

### v1

@@ 3 Main Claims
    $\operatorname{tr}^{S^{c}}|\phi_{1}\rangle\langle\phi_{1}|$ $\approx\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}$ (61a)
    $\operatorname{tr}^{S^{c}}|\phi_{1}\rangle\langle\phi_{2}|$ $\approx 0\,.$ (61b)
@@
    $\hat{\rho}_{\Psi_{t}}^{S}=\operatorname{tr}^{S^{c}}|\Psi_{t}\rangle\langle\Psi_{t}|\approx\hat{\rho}_{\mathrm{gG}}^{S}~~(\text{respectively, }\hat{\rho}_{\mathrm{gc}}^{S})\,.$ (62)
@@ 4.1 Comparison to Maximum Entropy Principle
First, the maximization problem leads to a representation of subjective probability (or the knowledge of an observer), but if statistical mechanics were limited to subjective probability then it would not be justified in most applications; in particular, it could not treat any events that take place in the absence of observers (e.g., the formation of stars before humans even existed). In contrast, there is nothing subjective about $\operatorname{tr}^{S^{c}}|\Psi\rangle\langle\Psi|$; that is, we have shown how $\hat{\rho}_{\mathrm{gc}}$ actually occurs in nature.
@@
The second issue above was mentioned by Balian [1, Sec. 4.1.2], who suggested that prescribing the expectation value of an observable $\hat{A}$ made sense when considering subsystems, of which there are many, and each may have a different value of $\hat{A}$. However, that seems like mixing up the description and the justification of $\hat{\rho}_{\mathrm{can}}$. If, as Balian’s reasoning suggests, $\hat{\rho}_{\mathrm{can}}$ can only apply to subsystems $S$, while only $\hat{\rho}_{\mathrm{mc}}$ can apply to full systems, then $\hat{\rho}^{S}_{\mathrm{can}}$ can only be justified as $\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{mc}}$, and the maximizer of $S_{\mathrm{vN}}$ for $S$ (under whichever constraints) would be irrelevant unless it agreed with $\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{mc}}$. That is, the reasoning apparently undercuts the use of $S_{\mathrm{vN}}$ for justifying $\hat{\rho}_{\mathrm{can}}$. For us, the consideration of subsystems also plays an important role, for example through the concept of local equivalence of density matrices, but the logic is different in our case, as we do not use subjective probability. $\diamond$
@@ 5 Derivation of Statement 1a: Generalized Gibbs Density Matrix
$=\Bigl[\hat{Q}_{k}^{S}\otimes\hat{I}^{S^{c}}+\hat{I}^{S}\otimes\hat{Q}_{k}^{S^{c}},\hat{Q}_{k^{\prime}}^{S}\otimes\hat{I}^{S^{c}}+\hat{I}^{S}\otimes\hat{Q}_{k^{\prime}}^{S^{c}}\Bigr]$ (68b)
$=\bigl[\hat{Q}_{k}^{S},\hat{Q}_{k^{\prime}}^{S}\bigr]\otimes\hat{I}^{S^{c}}+\hat{I}^{S}\otimes\bigl[\hat{Q}_{k}^{S^{c}},\hat{Q}_{k^{\prime}}^{S^{c}}\bigr]\,$, (68c)
@@
$\bigl[\hat{Q}_{k}^{S},\hat{Q}_{k^{\prime}}^{S}\bigr]=0=\bigl[\hat{Q}_{k}^{S^{c}},\hat{Q}_{k^{\prime}}^{S^{c}}\bigr]\,.$ (69)
For $X=S,S^{c}$ let $\{\phi_{j}^{X}:j\in\mathscr{J}^{X}\}$ be an ONB of $\mathscr{H}^{X}$ that is a joint eigenbasis of all $\hat{Q}_{k}^{X}$, and let $q_{kj}^{X}$ be the corresponding eigenvalue. With the notation $\mathscr{J}:=\mathscr{J}^{S}\times\mathscr{J}^{S^{c}}$ and $\phi_{jj^{\prime}}:=\phi^{S}_{j}\otimes\phi_{j^{\prime}}^{S^{c}}$, it follows that $\{\phi_{jj^{\prime}}:(j,j^{\prime})\in\mathscr{J}\}$ is an ONB of $\mathscr{H}$ that diagonalizes (among others) $\hat{Q}_{1},\ldots,\hat{Q}_{K}$. Let
@@
$=\Bigl\{(j,j^{\prime})\in\mathscr{J}\>:\>\forall k:q_{kj}^{S}+q_{kj^{\prime}}^{S^{c}}\in[Q_{k}-\Delta Q_{k},Q_{k}]\Bigr\}\,$, (70b)
@@
$\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}$ $=\frac{1}{d_{\mathrm{gmc}}}\sum_{(j,j^{\prime})\in\mathscr{J}_{\mathrm{gmc}}}\operatorname{tr}^{S^{c}}|\phi_{jj^{\prime}}\rangle\langle\phi_{jj^{\prime}}|$ (72a)
@@
$=\sum_{j\in\mathscr{J}^{S}}\frac{\#\{j^{\prime}\in\mathscr{J}^{S^{c}}:(j,j^{\prime})\in\mathscr{J}_{\mathrm{gmc}}\}}{d_{\mathrm{gmc}}}|\phi_{j}^{S}\rangle\langle\phi_{j}^{S}|$ (72c)
$=\sum_{j\in\mathscr{J}^{S}}\frac{\#\bigl\{j^{\prime}\in\mathscr{J}^{S^{c}}:\forall k:q_{kj^{\prime}}^{S^{c}}\in[Q_{k}-q_{kj}^{S}-\Delta Q_{k},Q_{k}-q_{kj}^{S}]\bigr\}}{d_{\mathrm{gmc}}}|\phi_{j}^{S}\rangle\langle\phi_{j}^{S}|$ (72d)
$=\sum_{j\in\mathscr{J}^{S}}\frac{\dim\mathscr{H}_{j}^{S^{c}}}{d_{\mathrm{gmc}}}|\phi_{j}^{S}\rangle\langle\phi_{j}^{S}|\,$, (72e)
@@
$\mathscr{H}_{j}^{S^{c}}:=\prod_{k=1}^{K}\mathds{1}_{[Q_{k}-q_{kj}^{S}-\Delta Q_{k},Q_{k}-q_{kj}^{S}]}(\hat{Q}_{k}^{S^{c}})\mathscr{H}^{S^{c}}.$ (73)
The Boltzmann entropy of the macro state of $S^{c}$ defined by the values $q_{k}^{S^{c}}$ with tolerance $\Delta Q_{k}$ is defined by [18, 17]
$S(q^{S^{c}}_{1},\ldots,q^{S^{c}}_{K}):=k_{\mathrm{B}}\log\dim\Biggl(\prod_{k=1}^{K}\mathds{1}_{[q^{S^{c}}_{k}-\Delta Q_{k},q^{S^{c}}_{k}]}(\hat{Q}_{k}^{S^{c}})\mathscr{H}^{S^{c}}\Biggr)$ (74)
with Boltzmann’s constant $k_{\mathrm{B}}$ and thus
$\dim\mathscr{H}_{j}^{S^{c}}=\exp\Biggl(\frac{1}{k_{\mathrm{B}}}S\Bigl(Q_{1}-q_{1j}^{S},\ldots,Q_{K}-q_{Kj}^{S}\Bigr)\Biggr).$ (75)
We now approximate the $S$ function in (75) by its Taylor expansion around $(Q_{1},\ldots,Q_{K})$. If the $\hat{Q}_{k},\hat{Q}_{k}^{S}$ are positive operators, this can be justified as follows: Since $S$ is small, $S^{c}$ is large compared to $S$, and positive extensive observables should be dominated by the contribution from $S^{c}$, so that, for most $(j,j^{\prime})\in\mathscr{J}_{\mathrm{gmc}}$ and all $k$,
$q_{kj}^{S}\ll q_{kj^{\prime}}^{S^{c}}\approx q_{kj}^{S}+q_{kj^{\prime}}^{S^{c}}\approx Q_{k}\,.$ (76)
@@
$S(Q_{1}-q_{1j}^{S},\dots,Q_{K}-q_{Kj}^{S})\approx S(Q_{1},\dots,Q_{K})-\sum_{k=1}^{K}\frac{\partial S}{\partial q^{S^{c}}_{k}}(Q_{1},\ldots,Q_{K})\,q_{kj}^{S}.$ (77)
@@
$\dim\mathscr{H}_{j}^{S^{c}}\approx\exp\bigl(S(Q_{1},\ldots,Q_{K})/k_{\mathrm{B}}\bigr)\>\exp\left(\sum_{k=1}^{K}\lambda_{k}q_{kj}^{S}\right)$ (78)
@@
$\lambda_{k}:=-\frac{1}{k_{\mathrm{B}}}\frac{\partial S}{\partial q_{k}^{S^{c}}}(Q_{1},\ldots,Q_{K})\,.$ (79)
(In particular, when this approximation is valid, the dimensions will be large.)
@@
$\operatorname{tr}^{S^{c}}\hat{\rho}_{\mathrm{gmc}}\approx\frac{1}{Z_{\mathrm{gG}}}\exp\left(\sum_{k=1}^{K}\lambda_{k}\hat{Q}^{S}_{k}\right),$ (80)
@@
It remains to verify that $\lambda_{k}$ given by (79) agrees (approximately) with $\lambda_{k}$ solving (10). Starting from (79), subdividing the available volume $\Lambda$ into many small regions $S_{\alpha}$ and using extensivity, we find that
$\operatorname{tr}(\hat{\rho}_{\mathrm{gG}}\,\hat{Q}_{k})$ $=\sum_{\alpha}\operatorname{tr}\Bigl[\hat{\rho}_{\mathrm{gG}}\,(\hat{Q}_{k}^{S_{\alpha}}\otimes\hat{I}^{S_{\alpha}^{c}})\Bigr]$ (81a)
$=\sum_{\alpha}\operatorname{tr}\Bigl[\operatorname{tr}^{S_{\alpha}^{c}}(\hat{\rho}_{\mathrm{gG}})\,\hat{Q}_{k}^{S_{\alpha}}\Bigr]$ (81b)
$\approx\sum_{\alpha}\operatorname{tr}\Bigl[\operatorname{tr}^{S_{\alpha}^{c}}(\hat{\rho}_{\mathrm{gmc}})\,\hat{Q}_{k}^{S_{\alpha}}\Bigr]$ (81c)
$=\sum_{\alpha}\operatorname{tr}\Bigl[\hat{\rho}_{\mathrm{gmc}}\,(\hat{Q}_{k}^{S_{\alpha}}\otimes\hat{I}^{S_{\alpha}^{c}})\Bigr]$ (81d)
@@
This completes the derivation of Statement 1a.

### v2

@@ 3 Main Claims
    $\operatorname{tr}^{\overline{S}}|\phi_{1}\rangle\langle\phi_{1}|$ $\approx\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}$ (65a)
    $\operatorname{tr}^{\overline{S}}|\phi_{1}\rangle\langle\phi_{2}|$ $\approx 0\,.$ (65b)
@@
    $\hat{\rho}_{\Psi_{t}}^{S}=\operatorname{tr}^{\overline{S}}|\Psi_{t}\rangle\langle\Psi_{t}|\approx\hat{\rho}_{\mathrm{gG}}^{S}~~(\text{respectively, }\hat{\rho}_{\mathrm{gc}}^{S})\,.$ (66)
@@ 4.1 Comparison to Maximum Entropy Principle
First, the maximization problem leads to a representation of subjective probability (or the knowledge of an observer), but if statistical mechanics were limited to subjective probability then it would not be justified in most applications; in particular, it could not treat any events that take place in the absence of observers (e.g., the formation of stars before humans even existed). In contrast, there is nothing subjective about $\operatorname{tr}^{\overline{S}}|\Psi\rangle\langle\Psi|$; that is, we have shown how $\hat{\rho}_{\mathrm{gc}}$ actually occurs in nature.
@@
The second issue above was mentioned by Balian [1, Sec. 4.1.2], who suggested that prescribing the expectation value of an observable $\hat{A}$ made sense when considering subsystems, of which there are many, and each may have a different value of $\hat{A}$. However, that seems like mixing up the description and the justification of $\hat{\rho}_{\mathrm{can}}$. If, as Balian’s reasoning suggests, $\hat{\rho}_{\mathrm{can}}$ can only apply to subsystems $S$, while only $\hat{\rho}_{\mathrm{mc}}$ can apply to full systems, then $\hat{\rho}^{S}_{\mathrm{can}}$ can only be justified as $\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{mc}}$, and the maximizer of $S_{\mathrm{vN}}$ for $S$ (under whichever constraints) would be irrelevant unless it agreed with $\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{mc}}$. That is, the reasoning apparently undercuts the use of $S_{\mathrm{vN}}$ for justifying $\hat{\rho}_{\mathrm{can}}$. For us, the consideration of subsystems also plays an important role, for example through the concept of local equivalence of density matrices, but the logic is different in our case, as we do not use subjective probability. $\diamond$
@@ 5 Derivation of Statement 1a: Generalized Gibbs Density Matrix
$=\Bigl[\hat{Q}_{k}^{S}\otimes\hat{I}^{\overline{S}}+\hat{I}^{S}\otimes\hat{Q}_{k}^{\overline{S}},\hat{Q}_{k^{\prime}}^{S}\otimes\hat{I}^{\overline{S}}+\hat{I}^{S}\otimes\hat{Q}_{k^{\prime}}^{\overline{S}}\Bigr]$ (72b)
$=\bigl[\hat{Q}_{k}^{S},\hat{Q}_{k^{\prime}}^{S}\bigr]\otimes\hat{I}^{\overline{S}}+\hat{I}^{S}\otimes\bigl[\hat{Q}_{k}^{\overline{S}},\hat{Q}_{k^{\prime}}^{\overline{S}}\bigr]\,$, (72c)
@@
$\bigl[\hat{Q}_{k}^{S},\hat{Q}_{k^{\prime}}^{S}\bigr]=0=\bigl[\hat{Q}_{k}^{\overline{S}},\hat{Q}_{k^{\prime}}^{\overline{S}}\bigr]\,.$ (73)
For $X=S,\overline{S}$ let $\{\phi_{j}^{X}:j\in\mathscr{J}^{X}\}$ be an ONB of $\mathscr{H}^{X}$ that is a joint eigenbasis of all $\hat{Q}_{k}^{X}$, and let $q_{kj}^{X}$ be the corresponding eigenvalue. With the notation $\mathscr{J}:=\mathscr{J}^{S}\times\mathscr{J}^{\overline{S}}$ and $\phi_{jj^{\prime}}:=\phi^{S}_{j}\otimes\phi_{j^{\prime}}^{\overline{S}}$, it follows that $\{\phi_{jj^{\prime}}:(j,j^{\prime})\in\mathscr{J}\}$ is an ONB of $\mathscr{H}$ that diagonalizes (among others) $\hat{Q}_{1},\ldots,\hat{Q}_{K}$. Let
@@
$=\Bigl\{(j,j^{\prime})\in\mathscr{J}\>:\>\forall k:q_{kj}^{S}+q_{kj^{\prime}}^{\overline{S}}\in[Q_{k}-\Delta Q_{k},Q_{k}]\Bigr\}\,$, (74b)
@@
$\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}$ $=\frac{1}{d_{\mathrm{gmc}}}\sum_{(j,j^{\prime})\in\mathscr{J}_{\mathrm{gmc}}}\operatorname{tr}^{\overline{S}}|\phi_{jj^{\prime}}\rangle\langle\phi_{jj^{\prime}}|$ (76a)
@@
$=\sum_{j\in\mathscr{J}^{S}}\frac{\#\{j^{\prime}\in\mathscr{J}^{\overline{S}}:(j,j^{\prime})\in\mathscr{J}_{\mathrm{gmc}}\}}{d_{\mathrm{gmc}}}|\phi_{j}^{S}\rangle\langle\phi_{j}^{S}|$ (76c)
$=\sum_{j\in\mathscr{J}^{S}}\frac{\#\bigl\{j^{\prime}\in\mathscr{J}^{\overline{S}}:\forall k:q_{kj^{\prime}}^{\overline{S}}\in[Q_{k}-q_{kj}^{S}-\Delta Q_{k},Q_{k}-q_{kj}^{S}]\bigr\}}{d_{\mathrm{gmc}}}|\phi_{j}^{S}\rangle\langle\phi_{j}^{S}|$ (76d)
$=\sum_{j\in\mathscr{J}^{S}}\frac{\dim\mathscr{H}_{j}^{\overline{S}}}{d_{\mathrm{gmc}}}|\phi_{j}^{S}\rangle\langle\phi_{j}^{S}|\,$, (76e)
@@
$\mathscr{H}_{j}^{\overline{S}}:=\prod_{k=1}^{K}\mathds{1}_{[Q_{k}-q_{kj}^{S}-\Delta Q_{k},Q_{k}-q_{kj}^{S}]}(\hat{Q}_{k}^{\overline{S}})\mathscr{H}^{\overline{S}}.$ (77)
The Boltzmann entropy of the macro state of $\overline{S}$ defined by the values $q_{k}^{\overline{S}}$ with tolerance $\Delta Q_{k}$ is defined by [21, 20]
$S(q^{\overline{S}}_{1},\ldots,q^{\overline{S}}_{K}):=k_{\mathrm{B}}\log\dim\Biggl(\prod_{k=1}^{K}\mathds{1}_{[q^{\overline{S}}_{k}-\Delta Q_{k},q^{\overline{S}}_{k}]}(\hat{Q}_{k}^{\overline{S}})\mathscr{H}^{\overline{S}}\Biggr)\,$, (78)
so that
$\dim\mathscr{H}_{j}^{\overline{S}}=\exp\Biggl(\frac{1}{k_{\mathrm{B}}}S\Bigl(Q_{1}-q_{1j}^{S},\ldots,Q_{K}-q_{Kj}^{S}\Bigr)\Biggr).$ (79)
We now approximate the $S$ function in (79) by its Taylor expansion around $(Q_{1},\ldots,Q_{K})$. If the $\hat{Q}_{k},\hat{Q}_{k}^{S}$ are positive operators, this can be justified as follows: Since $S$ is small, $\overline{S}$ is large compared to $S$, and positive extensive observables should be dominated by the contribution from $\overline{S}$, so that, for most $(j,j^{\prime})\in\mathscr{J}_{\mathrm{gmc}}$ and all $k$,
$q_{kj}^{S}\ll q_{kj^{\prime}}^{\overline{S}}\approx q_{kj}^{S}+q_{kj^{\prime}}^{\overline{S}}\approx Q_{k}\,.$ (80)
@@
$S(Q_{1}-q_{1j}^{S},\dots,Q_{K}-q_{Kj}^{S})\approx S(Q_{1},\dots,Q_{K})-\sum_{k=1}^{K}\frac{\partial S}{\partial q^{\overline{S}}_{k}}(Q_{1},\ldots,Q_{K})\,q_{kj}^{S}.$ (81)
@@
$\dim\mathscr{H}_{j}^{\overline{S}}\approx\exp\bigl(S(Q_{1},\ldots,Q_{K})/k_{\mathrm{B}}\bigr)\>\exp\left(\sum_{k=1}^{K}\lambda_{k}q_{kj}^{S}\right)$ (82)
@@
$\lambda_{k}:=-\frac{1}{k_{\mathrm{B}}}\frac{\partial S}{\partial q_{k}^{\overline{S}}}(Q_{1},\ldots,Q_{K})\,.$ (83)
@@
$\operatorname{tr}^{\overline{S}}\hat{\rho}_{\mathrm{gmc}}\approx\frac{1}{Z_{\mathrm{gG}}}\exp\left(\sum_{k=1}^{K}\lambda_{k}\hat{Q}^{S}_{k}\right),$ (84)
@@
For the next steps, we need the following mathematical fact, proved in Appendix D.
Proposition 1.
Let $\mathscr{H}^{a},\mathscr{H}^{b},\mathscr{H}^{c}$ be finite-dimensional Hilbert spaces. If for some operators
$\hat{Q}^{a}\otimes\hat{I}^{b}\otimes\hat{I}^{c}+\hat{I}^{a}\otimes\hat{Q}^{bc}=\hat{Q}^{ab}\otimes\hat{I}^{c}+\hat{I}^{a}\otimes\hat{I}^{b}\otimes\hat{Q}^{c}\,$, (85)
then there exists a unique operator $\hat{Q}^{b}$ such that
$\hat{Q}^{bc}$ $=\hat{Q}^{b}\otimes\hat{I}^{c}+\hat{I}^{b}\otimes\hat{Q}^{c}$ (86a)
$\hat{Q}^{ab}$ $=\hat{Q}^{a}\otimes\hat{I}^{b}+\hat{I}^{a}\otimes\hat{Q}^{b}\,.$ (86b)
As a consequence, if $S_{1}\subseteq S_{2}\subseteq\Lambda$, then from (13) for $S_{1}$ and $S_{2}$ we obtain that
$\hat{Q}_{k}$ $\approx\hat{Q}_{k}^{S_{1}}\otimes\hat{I}^{S_{2}\setminus S_{1}}\otimes\hat{I}^{\overline{S}_{2}}$
$+\hat{I}^{S_{1}}\otimes\hat{Q}_{k}^{S_{2}\setminus S_{1}}\otimes\hat{I}^{\overline{S}_{2}}$
$+\hat{I}^{S_{1}}\otimes\hat{I}^{S_{2}\setminus S_{1}}\otimes\hat{Q}_{k}^{\overline{S}_{2}}\,$, (87)
and (by repeated application of Proposition 1) similarly for any finite partition $\Lambda=\cup_{\alpha}S_{\alpha}$ ($S_{\alpha}\cap S_{\alpha^{\prime}}=\emptyset$ for $\alpha\neq\alpha^{\prime}$):
$\hat{Q}_{k}\approx\sum_{\alpha}\hat{Q}_{k}^{S_{\alpha}}\otimes\hat{I}^{\overline{S}_{\alpha}}\,.$ (88)
For the derivation of Statement 1a, it remains to verify that $\lambda_{k}$ given by (83) agrees (approximately) with $\lambda_{k}$ solving (10). Starting from (83), subdividing the available volume $\Lambda$ into many small regions $S_{\alpha}$ and using (88), we find that
$\operatorname{tr}(\hat{\rho}_{\mathrm{gG}}\,\hat{Q}_{k})$ $\approx\sum_{\alpha}\operatorname{tr}\Bigl[\hat{\rho}_{\mathrm{gG}}\,(\hat{Q}_{k}^{S_{\alpha}}\otimes\hat{I}^{\overline{S}_{\alpha}})\Bigr]$ (89a)
$=\sum_{\alpha}\operatorname{tr}\Bigl[\operatorname{tr}^{\overline{S}_{\alpha}}(\hat{\rho}_{\mathrm{gG}})\,\hat{Q}_{k}^{S_{\alpha}}\Bigr]$ (89b)
$\approx\sum_{\alpha}\operatorname{tr}\Bigl[\operatorname{tr}^{\overline{S}_{\alpha}}(\hat{\rho}_{\mathrm{gmc}})\,\hat{Q}_{k}^{S_{\alpha}}\Bigr]$ (89c)
$=\sum_{\alpha}\operatorname{tr}\Bigl[\hat{\rho}_{\mathrm{gmc}}\,(\hat{Q}_{k}^{S_{\alpha}}\otimes\hat{I}^{\overline{S}_{\alpha}})\Bigr]$ (89d)
@@
Thus, the $\lambda_{k}$ given by (83) satisfy (10). This completes the derivation of Statement 1a.